OpenStudy (anonymous):

Which expression is equivalent to 14a + 8? 14(a + 4) 7(7a + 1) 2(7a + 4) 14(a + 1)

OpenStudy (anonymous):

omg i have the same dong that you have ugh i miss mine hes on vacation ok to the question 14a + 8 use distributive property backwards :)

OpenStudy (anonymous):

dog*

OpenStudy (ryanchick):

14 times 8

OpenStudy (anonymous):

What does 14(a + 4) = ?

OpenStudy (ryanchick):

14 times 4

OpenStudy (anonymous):

-,- if I knew that I would have gotten the answer idk wat a is

OpenStudy (ryanchick):

times

OpenStudy (anonymous):

Are you familiar with the acronym FOIL ?

OpenStudy (anonymous):

do distributive property on all the questions until u get 14a + 8

OpenStudy (anonymous):

I think it is A though

OpenStudy (anonymous):

no that's 112

OpenStudy (anonymous):

14 x 4 isnt 8 :(

OpenStudy (anonymous):

Don't guess. Learn to understand how to work the problem so you can answer similar problems by yourself and get them correct

OpenStudy (anonymous):

14(a + 4) = 14*a+ 14*4 Does 14*4=8?

OpenStudy (anonymous):

try the same process here 7(7a + 1) does the answer match 14a + 8? if not, then try the next one

OpenStudy (ryanchick):

some of them yuo add and times and others you time you hafe to times in all of them

OpenStudy (ryanchick):

here are your answers 1.would be 56

OpenStudy (ryanchick):

next one 22

OpenStudy (ryanchick):

last one 14

OpenStudy (ryanchick):

hope those are helpfull

OpenStudy (anonymous):

the important thing here is that i tried to help...and not give out answers